# Tweezer interferometry with NOON states

### v1

@@ IV Hamiltonian formulation
The dimer Hamiltonian can be written using generators of spin-rotations. One starts with standard Bose-Hubbard version, namely,
@@
In a quantum context $(S_{x},S_{y},S_{z})$ are operators, while in a semiclassical perspective these are “phasespace” coordinates of the so-called Bloch sphere, with ${S_{x}^{2}+S_{y}^{2}+S_{z}^{2}\approx(N/2)^{2}}$. The contour lines ${H(S_{x},S_{y},S_{z})=E}$ are the skeleton on which the quantum states are constructed. In a WKB picture (see Appendix A) each contour line supports an eigenstate, and the area spacing between contour lines equals the Planck cell area $h\approx 2\pi$. See the upper panel of Fig. 3 for a demonstration. Any state of the system can be represented on the Bloch sphere using e.g. Wigner or Husimi representation.
@@
For $u{=}0$ the energy landscape has one minimum (at the East pole ${S_{x}=N/2}$), and one maximum (at the West pole ${S_{x}=-N/2}$). Considering an unbiased system with $\Delta=0$, as $J$ is decreased, a bifurcation appears at the East pole once $u$ becomes larger than unity. At this point, a double well structure is formed, as illustrated in the upper panel of Fig. 3. The minima are located at $S_{x}=u^{-1}(N/2)$ with $S_{z}=\pm\sqrt{1-u^{-2}}(N/2)$, while $S_{y}=0$. In particular, for $J=0$ the minima are at the North and South poles.
@@
$J_{c}=\left[(NU)^{2/3}-\Delta^{2/3}\right]^{3/2}$ (15)
For $J\lesssim J_{c}$ the potential becomes an asymmetric double-well structure with a secondary minimum that is born at the vicinity of the East pole. The bifurcation for ${\Delta\neq 0}$ is of a saddle-node type, instead of the ${\Delta=0}$ pitchfork bifurcation that generates a symmetric double well.
Figure 4: Dimer frequencies: Upper panel: The dependence of $\omega(E)$ on the energy is extracted (black dots) from the difference $E_{\nu{+}1}-E_{\nu}$ for $N=500$ particles with $UN=1$ and $J=0.4$. Here $\Delta=0$ and therefore the lower part of the spectrum consists of doublets, which is reflected by the appearance of an $\omega\sim 0$ branch. The exact and approximated analytical results Eq. (A) and Eq. (39) are represented by green and red lines. Lower panel: The characteristic frequency near the separatrix ($\omega_{x}$) is plotted as a function of $J$. The Black line is from the calculation of energy differences; the blue line is $\omega_{J}$ of Eq. (16); and the red line is the analytical estimate Eq. (18). The red asterisk is Eq. (17).
The stationary states of the Hamiltonian have energies $E_{\nu}$. A superposition of two consecutive non-degenerate eigenstates with oscillates with a frequency $\omega(E)$ that is determined by the level spacing $E_{\nu{+}1}-E_{\nu}$. Fig. 4 demonstrates the energy dependence of $\omega(E)$, which can be analytically determined using a WKB procedure, see Appendix A. The characteristic frequency ${\omega_{0}\equiv NU}$ serves to fix the time units in our simulations. For ${J>J_{c}}$ and ${\Delta=0}$ the frequency at the minimum energy is
@@
Thus, the frequency $\omega_{x}$ of Eq. (18) becomes zero in the classical limit (${N=\infty}$), but away from the bifurcation it is finite and of order $\omega_{J}$ for any realistic finite $N$. This observation is crucial for the design of dynamical splitting. In the non-adiabatic scheme, the time to get an approximate cat state is $\pi/\omega_{x}$. The outcome is illustrated on the right side of the lower panel of Fig. 3. On the left side of the same panel we display the outcome of an adiabatic splitting process. In the latter case, a finite level spacing is an essential condition. The bottleneck of the adiabatic splitting process is during the bifurcation of the EastPole, which happens when the seperatrix is born, where Eq. (17) rather than Eq. (18) determines the level spacing. We discuss and analyze the implied adiabatic condition in Section VI.
Figure 5: Adiabatic evolution. Illustration of representative simulations with $N{=}10$ particles. The evolving state is projected on the adiabatic states. The probabilities are color-coded. If the probability $p_{\nu}$ is negligibly small, the color is gray. Red color indicates ${p\approx 1}$. The accumulated (added) relative phase at $t{=}0$ is $\varphi{=}0$ (upper) and $N\varphi{=}\pi/3$ (middle) and $N\varphi{=}\pi$ (lower). The units of time are chosen such that ${NU=1}$. The sweep rate is $\dot{J}=10^{-4}$ and $\dot{\Delta}=10^{-4}$, while $J(\text{max}){=}1.1$, and $\Delta(\text{max})=U/2$.
@@ V Simulations
where $\left|E_{\nu}(t)\right\rangle$ are the adiabatic levels, namely, the instantaneous eigen-energies of the time dependent Hamiltonian $H(t)$. At the end of the protocol $J{=}0$ while ${\Delta\neq 0}$, and accordingly the eigenstates of the non-degenerated Hamiltonian $H(t)$ are the occupation states ${|N-n,n\rangle}$. It follows that at the end of the simulation the probabilities $p_{\nu}$, up to indexing, are the probabilities $P(n)$ that we want to calculate. As explained in the previous paragraph, in an ideal adiabatic protocol each run of the experiment ends with either ${n=0}$ or ${n=N}$. In practice, due to lack of strict adiabaticity, or due to other disturbances we have non-zero probability to get other values of $n$.
In Fig. 6 we show the dependence of $\left\langle n\right\rangle$ on $\varphi$ for representative sweep rates without or with rejection of the “bad” runs. The “bad” runs are those for which the outcome is neither ${n=0}$ nor ${n=N}$, indicating an undesired non-adiabatic transition. Due to this rejection we get a filtered estimate $\left\langle n\right\rangle_{f}$ instead of $\left\langle n\right\rangle$, namely,
@@ VI The optimized adiabatic process
The upper panel of Fig. 7 shows the non-adiabatic leakage to higher levels. The gross dependence of the expectation value $\left\langle n_{\text{ex}}\right\rangle$ on the sweep rate is similar for the semiclassical and quantum calculations. Our interest is in the quantum adiabatic regime, where we want to maintain ${\left\langle n_{\text{ex}}\right\rangle\ll 1}$. As can be seen in the figure, for a given bound that we want to impose on $\left\langle n_{\text{ex}}\right\rangle$, the maximum allowed sweep rate depends on $N$.
The quality of the phase estimation depends on the adiabaticity, and hence on our ability to allocate sufficient time for the splitting ($\Delta=0$) and branching ($\Delta\neq 0$) sweeps, during which $J$ is decreased to zero. On the other hand, practical considerations give preference to shorter sequences. This motivates us to suggest an optimized adiabatic process, where the sweep rate is continuously modified.
@@
where $C$ is a numerical constant. Representative plots of $F(J)$, and the dependence of $\max{F(J)}$ and $T$ of $N$ are provided in Fig. 8. Recall that we fix the time units such that ${NU=1}$. Based on these results we conclude that ${\alpha\sim 0}$. Thanks to the optimization, the required time for an adiabatic sweep is shortened, albeit with the same scaling with respect to $N$, and for the numerical prefactor we get ${C\approx 0.07}$. We note that the semiclassical analysis suggests ${N\log(N)}$ dependence, but we cannot resolve the logarithmic correction numerically.
@@
Whether we get splitting or branching depends on $\Delta/U$. An important question to address is what is the tolerance of each of the processes to small deviations in setting $\Delta/U$. This question is particularly important considering the feasibility of our suggested protocol. The answer is provided in Fig. 9, where we plot the mean occupation $\left\langle n\right\rangle$ at the end of the sweep of $J$ down to zero. This quantity can be used as an indicator to whether we get a cat-state with $\left\langle n\right\rangle=N/2$ or branching to either $\left\langle n\right\rangle=0$ or $\left\langle n\right\rangle=N$. We observe that for increased $N$, not only the tolerance does not deteriorate, but it even improves. Thus, in practice, one needs to tune the bias $\Delta$ to zero with an accuracy better than $0.01\times U$.
The merging process.– The protocol of Fig. 1 includes a merging stage that constitutes a reversed version of the splitting. It maps the EvenCat and OddCat states to $|N,0\rangle$ and $|N{-}1,1\rangle$ respectively. We have verified (not displayed) that the same adiabatic condition applies.
@@
VII Crossover from adiabatic to diabatic scenario
[…]
In the adiabtic range, the leakage $\left\langle n_{\text{ex}}\right\rangle$ to higher levels is zero, because the final state is a superposition of the $S_{z}=-N/2$ and $S_{z}=N/2$ states for which $n_{\text{ex}}=0$. In the diabatic range, the sudden approximation implies that the state stays at $S_{x}=N/2$, therefore $n_{\text{ex}}\sim N/2$.
VIII Feasibility of the proposed protocol
@@
There are two parameters of each of the trapping potentials: $V_{0}$ and $\sigma$. To estimate the number of atoms, we take atom density of $\rho=10^{14}\,\text{cm}^{-3}$, which is typical for a BEC of ultracold atoms. The interaction energy scale is given by $U=4\pi\hbar^{2}\rho a$, where $a$ is the s-wave scattering length. The sweeps of J should be done on a scale of $NU$, and therefore, we require $NU\lesssim\Omega_{r}$. For 87Rb, the scattering length is on the order of $100a_{0}$, with $a_{0}$ being the Bohr radius. This yields an interaction energy scale on the order of $2\pi\times 1$ kHz, which means that for $N\sim 10-100$, we get that $J$ needs to be a few tens of kHz. This translates into a stringent requirement for a tightly confined trap. However, the size of the condensate in a tightly confined trap is very small, which together with a given density means a small number of atoms.
@@
To solve these apparently conflicting requirements, one needs to allow tunability of the interaction energy. This can be achieved by working near a magnetic Feshbach resonance. We thus suggest working with a condensate of 39K atoms at the state $\left|f=1,m_{f}=1\right\rangle$, which have a broad resonance near $B_{0}=403.4$G [28, 29]. Importantly, the scattering length can be reduced all the way to the point where it crosses zero, around $B_{ZC}=350$G [30]. Since adiabaticity requires that the sweep times are much larger than $U^{-1}$, we need to choose $a$ such that the interaction is not too small.
@@
As an example, let us fix $a=5a_{0}$ and choose a waist of $\sigma=10\mu$m, which does not require an objective objective with particularly high numerical aperture. With these numbers, we need to set the trap depth to $V_{0}=1.28\times 10^{-27}$ Joule, which amounts to approximately $93\mu$K. Assuming a trapping laser with a wavelength of $1064$nm, far off the atomic resonance, this potential depth translates into a modest laser power of around $100$mW [31].
@@
By the choice of parameters, we are working in the weak-interactions limit of the BEC [32], hence we shall use a Gaussian approximation for its shape. Thus, the number of atoms in the condensate is given by $N=\rho\pi^{3/2}w_{r}^{2}w_{z}$, with ${w_{r/z}=\sqrt{\hbar/(m\Omega_{r/z})}}$ being the ground state extent in each direction [32]. With our choice of parameters, we obtain $N\approx 50$, and $U/\hbar\approx 2\pi\times 86$Hz, and $J/\hbar\approx 2\pi\times 4300$Hz, and $\Omega_{r}\approx 2\pi\times 4478$Hz. The hierarchy of energy scales satisfy the required conditioned. Based on the results in Fig. 7 and Fig. 8, the sweep time of $J$ should be longer than $10\cdot 0.07(U/h)^{-1}\sim 8$ms. The most challenging aspect of the experiment is the requirement to stabilize the bias between the two potentials to within $0.01\times U$, which required a relative stability of $\sim 0.5\cdot 10^{-6}$ in the laser power. A plausible approach to generate a BEC of 50 atoms in a small trap is to load the trap from a large BEC that acts as a reservoir [33].
@@
IX Discussion
Several ways have been suggested how to prepare a NOON state. One way is to use the Schrodinger cat procedure, as in the experiment of Ref. [26], which requires the control and measurement of an internal degree of freedom. Another way is to use one-axis twisting, as described after Eq(90) of Ref. [2]. It seems that this latter method is sensitive to timing issues, and there may be complications that are related to parity measurement. Consequently, an optional way has been suggested [6] for the preparation of an approximate NOON state, namely, stretching a Gaussian cloud along a seperatrix, see Fig. 3b. The quality of the prepared cat state can be judged by looking at the final $P(n)$ distribution, see upper panel of Fig. 10. Note that for larger $N$ a deterioration is observed in the height of the $n{=}0$ and $n{=}N$ peaks. Ideally we would like to observe ${P(0)=P(N)=0.5}$.
In view of the above, our suggested procedure for phase-estimation is an adiabatic version of non-linear splitting process. Such method is robust, not sensitive to timing issues, and provides better interferometry for large $N$. See illustration of the outcome in the middle panel of Fig. 10b, where the $P(n)$ reflects an ideal NOON state.

### v2

@@ IV Hamiltonian formulation
The dimer Hamiltonian can be written using generators of spin-rotations. One starts with the standard Bose-Hubbard version, namely,
@@
In a quantum context $(S_{x},S_{y},S_{z})$ are operators of an $S=N/2$ spin entity, while in a semiclassical perspective these are “phasespace” coordinates of the so-called Bloch Sphere, with ${S_{x}^{2}+S_{y}^{2}+S_{z}^{2}=(S{+}1)S\approx(N/2)^{2}}$. The contour lines ${H(S_{x},S_{y},S_{z})=E}$ are the skeleton on which the quantum states are constructed. In a WKB picture (see Appendix A) each contour line supports an eigenstate, and the area spacing between contour lines equals the Planck cell area $h\approx 2\pi$. See the upper panel of Fig. 3 for a demonstration. Any state of the system can be represented on the Bloch sphere using e.g. Wigner or Husimi representation.
@@
For $u{=}0$ the energy landscape on the Bloch sphere has one minimum (at ${S_{x}=N/2}$), and one maximum (at ${S_{x}=-N/2}$). Considering an unbiased system with $\Delta=0$, as $J$ is decreased, a bifurcation appears at ${S_{x}=N/2}$ once $u$ becomes larger than unity. At this point, a double well structure is formed, as illustrated in the upper panel of Fig. 3. The minima are located at $S_{x}=u^{-1}(N/2)$ with $S_{z}=\pm\sqrt{1-u^{-2}}(N/2)$, while $S_{y}=0$. In particular, for $J=0$ the minima are at the North and South poles.
@@
$J_{c}=\left[(NU)^{2/3}-\Delta^{2/3}\right]^{3/2}\ \ .$ (17)
For $J\lesssim J_{c}$ the potential becomes an asymmetric double-well structure with a secondary minimum that is born at the vicinity of the ${S_{x}=N/2}$. The bifurcation for ${\Delta\neq 0}$ is of a saddle-node type, instead of the ${\Delta=0}$ pitchfork bifurcation that generates a symmetric double well.
Figure 4: Dimer frequencies: Upper panel: The dependence of $\omega(E)$ on the energy is extracted (black dots) from the difference $E_{\nu{+}1}-E_{\nu}$ for $N=500$ particles. The dimensionless energy and frequency correspond to choice of units such that ${NU=1}$, while $J=0.4$. Here $\Delta=0$ and therefore the lower part of the spectrum consists of doublets, which is reflected by the appearance of an $\omega\sim 0$ branch. The exact and approximated analytical results Eq. (35) and Eq. (41) are represented by green and red lines. Lower panel: The characteristic frequency near the separatrix ($\omega_{x}$) is plotted as a function of $J$. The Black line is from the calculation of energy differences; the blue line is $\omega_{J}$ of Eq. (18); and the red line is the analytical estimate Eq. (20). The red asterisk is Eq. (19).
The stationary states of the Hamiltonian have energies $E_{\nu}$. A superposition of two consecutive non-degenerate eigenstates oscillates with a frequency $\omega(E)$ that is determined by the level spacing $E_{\nu{+}1}-E_{\nu}$. Fig. 4 demonstrates the energy dependence of $\omega(E)$, which can be analytically determined using a WKB procedure, see Appendix A. The characteristic frequency ${\omega_{0}\equiv NU}$ serves to fix the time units in our simulations. For ${J>J_{c}}$ and ${\Delta=0}$ the frequency at the minimum energy is
@@
Thus, the frequency $\omega_{x}$ of Eq. (20) becomes zero in the classical limit (${N=\infty}$), but away from the bifurcation it is finite and of order $\omega_{J}$ for any realistic finite $N$. This observation is crucial for the design of dynamical splitting. In the non-adiabatic scheme, the time to get an approximate cat state is $\pi/\omega_{x}$. The outcome is illustrated on the right side of the lower panel of Fig. 3. On the left side of the same panel we display the outcome of an adiabatic splitting process. In the latter case, a finite level spacing is an essential condition. The bottleneck of the adiabatic splitting process is during the bifurcation at ${S_{x}=N/2}$, which happens when the separatrix is born, where Eq. (19) rather than Eq. (20) determines the level spacing. We discuss and analyze the implied adiabatic condition in Section VI.
Figure 5: Adiabatic evolution. Illustration of representative simulations with $N{=}10$ particles. The stages (a)-(e) of the protocol are indicated. The adiabatic levels $E_{\nu}$ are plotted as a function of time. The evolving state is projected on the adiabatic states. The probabilities are color-coded. If the probability $p_{\nu}$ is negligibly small, the color is gray. Red color indicates ${p\approx 1}$. The accumulated (added) relative phase at $t{=}0$ is $\varphi{=}0$ (upper) and $N\varphi{=}\pi/3$ (middle) and $N\varphi{=}\pi$ (lower). The units of time and energy are chosen such that ${NU=1}$. The sweep rate is $\dot{J}=10^{-4}$ and $\dot{\Delta}=10^{-4}$, while $J(\text{max}){=}1.1$, and $\Delta(\text{max})=U/2$.
@@ V Simulations
where $\left|E_{\nu}(t)\right\rangle$ are the adiabatic levels, namely, the instantaneous eigen-energies of the time dependent Hamiltonian $H(t)$. At the end of the protocol $J{=}0$ while ${\Delta\neq 0}$, and accordingly the eigenstates of the non-degenerated Hamiltonian $H(t)$ are the occupation states ${|N-n,n\rangle}$. It follows that at the end of the simulation the probabilities $p_{\nu}$, up to indexing, are the probabilities $P(n)$ that we want to calculate.
Idealized protocol.– As explained in the previous paragraph, in an ideal adiabatic protocol each run of the experiment ends with either ${n=0}$ or ${n=N}$. In practice, due to lack of strict adiabaticity, or due to other disturbances we have non-zero probability to get other values of $n$. This bring the idea that a rejection of bad measurements can improve the results. We discuss this option below. On top there might be errors that might not be as easily identified. We discuss briefly this point below as well.
Handling of explicit errors.– By explicit error we mean getting in some runs illegal values of $n$. In Fig. 6 we show the dependence of $\left\langle n\right\rangle$ on $\varphi$ for representative sweep rates without or with rejection of the “bad” runs. The “bad” runs are those for which the outcome is neither ${n=0}$ nor ${n=N}$, indicating an undesired non-adiabatic transition. Due to this rejection we get a filtered estimate $\left\langle n\right\rangle_{f}$ instead of $\left\langle n\right\rangle$, namely,
@@
Handling other systematic errors.– There are errors whose testing requires to go beyond the framework of the Bose-Hubbard modeling. Let us focus on one example. Let us assume that there is dephasing during the phase-accumulation stage. Such dephasing can occur e.g. due to the loss of even a single particle. Thus, instead of getting a coherent cat superposition ${\left|\varphi\right\rangle}$, we get a mixture of ${\left|Z\right\rangle\propto\left|+\right\rangle+\left|-\right\rangle}$ and ${\left|\bar{Z}\right\rangle\propto\left|+\right\rangle-\left|-\right\rangle}$. The outcome of the latter cannot be distinguished from the outcome of ${\left|\varphi=\pm\pi/2\right\rangle\propto\left|+\right\rangle\pm i\left|-\right\rangle}$ because in all these cases we get equal probabilities for $n{=}0$ and $n{=}N$. Accordingly, what we measure in practice is
$\frac{P(N)}{P(0)}\ \ \mapsto\ \ \frac{(1-p)P(N)+p}{(1-p)P(0)+p}$ (23)
where $p$ is the probability to have a dephasing events (or loss of particle) during phase accumulation. The good news here is that such error has a systematic effect on the statistics, and therefore, after extraction of $p$, can be compensated.
@@ VI The optimized adiabatic process
We are dealing with an adiabatic scheme, and therefore the crucial question is whether the required time for its implementation is realistic. The purpose of this section is to provide an estimate for the required time, and to see how it scales with $N$. Roughly speaking the variation of the control parameter should be slow if the occupied adiabatic levels are getting close, and can be faster if they are weakly coupled or well spaced apart.
The upper panel of Fig. 7 shows the non-adiabatic leakage to higher levels. The gross dependence of the expectation value $\left\langle n_{\text{ex}}\right\rangle$ on the sweep rate is similar for the semiclassical and quantum calculations. Appendix B provides further details regrading the crossover from the adiabatic to the diabatic scenario. Our interest is in the quantum adiabatic regime, where we want to maintain ${\left\langle n_{\text{ex}}\right\rangle\ll 1}$. As can be seen in the figure, for a given bound that we want to impose on $\left\langle n_{\text{ex}}\right\rangle$, the maximum allowed sweep rate depends on $N$.
The quality of the phase estimation depends on the adiabaticity, and hence on our ability to allocate sufficient time for the splitting ($\Delta=0$) and branching ($\Delta\neq 0$) sweeps, during which $J$ is decreased to zero. On the other hand, practical considerations give preference to shorter sequences. This motivates us to suggest an optimized adiabatic process, where the sweep rate is continuously modified. Similar idea has been considered in [28].
@@
where $C$ is a numerical constant. Representative plots of $F(J)$, and the dependence of $\max{F(J)}$ and $T$ of $N$ are provided in Fig. 8. Recall that we fix the time units such that ${NU=1}$. Based on these results we conclude that ${\alpha\sim 0}$. Thanks to the optimization, the required time for an adiabatic sweep is shortened, albeit with the same scaling with respect to $N$, and for the numerical prefactor we get ${C\approx 0.07}$. We note that the semiclassical analysis suggests ${N\ln(N)}$ dependence, but we cannot resolve the logarithmic correction numerically.
@@
Whether we get splitting or branching depends on $\Delta/U$. It is important to figure out what is the tolerance of each of the processes to small deviations in setting $\Delta/U$. This determines the feasibility of the suggested protocol. The answer is provided in Fig. 9, where we plot the mean occupation $\left\langle n\right\rangle$ at the end of the sweep of $J$ down to zero. This quantity can be used as an indicator to whether we get a cat-state with $\left\langle n\right\rangle=N/2$ or branching to either $\left\langle n\right\rangle=0$ or $\left\langle n\right\rangle=N$. We observe that for increased $N$, not only the tolerance does not deteriorate, but it even improves. Thus, in practice, one needs to tune the bias $\Delta$ to zero with an accuracy better than $0.01\times U$.
The merging process.– The protocol of Fig. 1 includes a merging stage that constitutes a reversed version of the splitting. It maps the EvenCat and OddCat states to the ground state $\left|X\right\rangle$ and to its one-particle excitation-state respectively. We have verified (not displayed) that the same adiabatic condition applies.
@@
VII Feasibility of the protocol
@@
There are two parameters of each of the trapping potentials: $V_{0}$ and $\sigma$. To estimate the number of atoms, we take atom density of $\rho=10^{14}\,\text{cm}^{-3}$, which is typical for a BEC of ultracold atoms. The interaction energy scale is given by $U=4\pi\hbar^{2}\rho|a|$, where $a$ is the s-wave scattering length. The sweeps of J should be done on a scale of $NU$, and therefore, we require $NU\lesssim\Omega_{r}$. For the gain in using a NOON state relative to a classical state to be significant, we aim to have $N\sim 10-100$ atoms.
@@
To fulfill all the requirements, one needs to allow tunability of the interaction energy. This can be achieved by working near a magnetic Feshbach resonance. We thus suggest working with a condensate of 39K atoms at the state $\left|f=1,m_{f}=1\right\rangle$, which have a broad resonance near $B_{0}=403.4$G [29, 30]. Importantly, the scattering length can be reduced all the way to the point where it crosses zero, around $B_{ZC}=350$G [31]. Since adiabaticity requires that the sweep times are much larger than $U^{-1}$, we need to choose $a$ such that the interaction is not too small.
@@
As an example, let us fix $a=-5a_{0}$ and choose a waist of $\sigma=10\mu$m, which does not require an objective with particularly high numerical aperture. With these numbers, we need to set the trap depth to $V_{0}=1.28\times 10^{-27}$ Joule, which amounts to approximately $93\mu$K. Assuming a trapping laser with a wavelength of $\lambda_{l}=1064$nm, far off the atomic resonance, this potential depth translates into a modest laser power of around $P_{0}=100$mW [32].
@@
By the choice of parameters, we are working in the weak-interactions limit of the BEC [33], hence we shall use a Gaussian approximation for its shape. Thus, the number of atoms in the condensate is given by $N=\rho\pi^{3/2}w_{r}^{2}w_{z}$, with ${w_{r/z}=\sqrt{\hbar/(m\Omega_{r/z})}}$ being the ground state extent in each direction [33]. With our choice of parameters, we obtain $N\approx 50$, and $U/\hbar\approx 2\pi\times 86$Hz, and $J/\hbar\approx 2\pi\times 4300$Hz, and $\Omega_{r}\approx 2\pi\times 4478$Hz. The hierarchy of energy scales satisfy the required conditions. For an attractive BEC to be stable against collapse, $N<N_{c}$ with $N_{c}\approx 0.57{(w_{r}^{2}w_{z})^{1/3}}/{|a|}$, see [34]. For our choice of parameters $N_{c}\approx 973$, therefore a BEC with $50$ atoms is stable. A plausible approach to generate a BEC of 50 atoms in a small trap is to load the trap from a large BEC that acts as a reservoir [35].
Based on the results in Fig. 7 and Fig. 8, the sweep time of $J$ should exceed $10\times 0.07(U/h)^{-1}\sim 8$ ms. The most challenging aspect of the experiment is the requirement to stabilize the bias between the two potentials to within $0.01\times U$, necessitating a relative stability of $\sim 0.5\cdot 10^{-6}$ in the laser power. This level of stability requires a high dynamic range for monitoring the beam’s power, achievable by combining multiple detectors with varying sensitivities. If we limit the shortest response time of the laser intensity locking circuit to $\tau=1$ ms, more than sufficient for executing our proposed protocol, the relative shot noise in the laser intensity is $\sqrt{{hc}/{(\lambda_{l}\tau P_{0})}}\approx 4\cdot 10^{-8}$, an order of magnitude below the required stability. Thus, we conclude that the necessary laser intensity stability is feasible.
@@
VIII Discussion
Several ways have been suggested how to prepare a NOON state. One way is to use the Schrodinger cat procedure, as in the experiment of Ref. [14], which requires the control and measurement of an internal degree of freedom. Other ways, that do not require an auxiliary degree of freedom, are e.g. one-axis twisting, as described after Eq(90) of Ref. [1]. Optionally, Ref.[5] has suggested to stretch a Gaussian cloud along a separatrix, see Fig. 3b. The quality of the prepared cat state can be judged by looking at the final $P(n)$ distribution, see upper panel of Fig. 10. Note that for larger $N$ a deterioration is observed in the height of the $n{=}0$ and $n{=}N$ peaks. Ideally we would like to observe ${P(0)=P(N)=0.5}$. Irrespective of that, the above methods are sensitive to timing issues, and there are complications that are related to the lack of experimental control over the exact number of particles, and associated difficulties in parity measurements.
In view of the above, our preferred procedure for phase-estimation is an adiabatic version of non-linear splitting process. Such method is robust, suitable for the tweezer interferometry setup, not sensitive to timing issues, and provides better interferometry for large $N$. See illustration of the outcome in the middle panel of Fig. 10b, where the $P(n)$ reflects an ideal NOON state.